OpenStudy (jade27p):

As part of an Earth Science project, Erin tracked the amount of rain fall for 4 weeks. She started tracking after finding some initial data on the year's rainfall to date: Week Rainfall 1 5.5 2 8 3 10.5 4 13 If the linear function representing the amount of rainfall each week is f(w) = 2.5w + 3, what does the 3 represent in this function? The amount of rainfall in week 0 The week that Erin collected data The amount of total rainfall The amount the rainfall increased each week

OpenStudy (jadeishere):

It'd be the initial rainfall, so, in other words, Week 0's rainfall total. For every number you can do this; (2.5 +3) = 5.5 2.5(2) + 3 = 8 2.5(3) + 3 = 10.5 2.5(4) + 3 = 13